1.     A car accelerates at a rate of 0.4ms-2 due to an unbalanced force of 500N.  Calculate the mass of the car.



2.     A boat of mass 2000kg is propelled forward due to an unbalanced force of 6500N.  Calculate the acceleration of the boat.




3.     If a 2kg block is pulled with a force of 10N and friction is 2N

a.     What is the size of the unbalanced force?

b.     What is the blocks acceleration?




4.     Two girls push a car of mass 1000 kg.  Each pushes with a force of 100 N and the force of friction is 120 N.  Calculate the acceleration of the car.



5.     A boat engine produces a  force of 10000 N and the friction and water resistance total 3500 N.  If the mass of the boat is 2000 kg, what will be its acceleration?















1
Expert's answer
2021-04-16T07:26:28-0400

1) "F=m\times a"

"500=m\times 0.40"

m=1250kg

2)"F=m\times a"

"2000\times a=6500"

"a=3.25\frac{m}{s^2}"

3)

a) Net force=10-2=8N

"Net Force=m\times a"

"8=2\times a"

"a=4\frac{m}{s^2}"

4) Net force F=200-120=80N

"Net force F=m\times a"

"80=100\times a"

"a=0.08\frac{m}{s^2}"

5) Net force=10000-350=9650N

"9650=2000\times a"

"a=4.825\frac{m}{s^2}"